OpenStudy (anonymous):

I need help bad!!! I need to find out and show work for 2x=x+2 and 2x+4=x+6

OpenStudy (anonymous):

2x=x+2 -x -x x=2 2x+4=x+6 x+4=6 x=2

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

2x=x+2 (2x/2)= (x+2)/2 x=2 2x+4=x+6 -4 -4 2x=x+2 x=2

OpenStudy (anonymous):

no problem:)

OpenStudy (anonymous):

lol i did /2 and you did -x. both work the same way

OpenStudy (anonymous):

thank u 4 the help!!!

OpenStudy (anonymous):

no prob